1. The stock of the McCall Corporation is currently trading at $42 per share. The stock’s volatility as measured by its standard deviation is 20%. If the strike (exercise) price for a certain set of options on McCall stock carry a strike price of $40, and the options run for 6 months (180 days), determine the Black-Scholes model values for:

N (d1), N (d2), N (- d1), and N (- d2). (Assume the risk-free rate is 10% and that the stock pays no dividends.)

2. Given the information in question one (1) above and your calculated values for N (d1), N (d2), N (- d1), and N (- d2), determine: a. The price (premium) one must pay for a call option (Ct) on McCall stock using the Black-Scholes model.

b. How much of the call’s value is composed of intrinsic value? How much of the call’s value is composed of a time premium? (Explain how you determined these values.)

3. Given the information in question one (1) above and your calculated values for N (d1), N (d2), N (- d1), and N (- d2), determine:

a. The price (premium) of a put option (Pt) for McCall stock using the BlackScholes model. (Use the put/call parity model to confirm your answer.)

b. How much of the put’s value (price) is composed of intrinsic value? How much of the put’s value is composed of a time premium? (Explain how you determined these values.)

These are the answers below, how do I get to them?

1. N (d1) = 0.7792 (where d1 = 0.7694) N (d2) = 0.7350 (where d2 = 0.6280) N (- d1) = 0.2208 N (- d2) = 0.2650

2. a. $4.76 (call price) b.

3. a. $0.81 (put price) b.
Answer #1

We use Black-Scholes Model to calculate the value of the call and put options.

The value of a call and put option are:

C = (S0 * N(d1)) - (Ke-rt * N(d2))

P = (K * e-rt)*N(-d2) - (S0)*N(-d1)

where :

S0 = current spot price

K = strike price

N(x) is the cumulative normal distribution function

r = risk-free interest rate

t is the time to maturity in years

d1 = (ln(S0 / K) + (r + σ2/2)*T) / σ√T

d2 = d1 - σ√T

σ = standard deviation of underlying stock returns

First, we calculate d1 and d2 as below :

  • ln(S0 / K) = ln(42 / 40). We input the same formula into Excel, i.e. =LN (42 / 40)
  • (r + σ2/2)*T = (0.10 + (0.202/2)*0.50
  • σ√T = 0.20 * √0.50

d1 = 0.7693

d2 = 0.6278

N(d1), N(-d1), N(d2),N(-d2) are calculated in Excel using the NORMSDIST function and inputting the value of d1 and d2 into the function.

N(d1) = 0.7791

N(d2) = 0.7349

N(-d1) = 0.2209

N(-d2) = 0.2651

Now, we calculate the values of the call and put options as below:

C = (S0 * N(d1))   - (Ke-rt * N(d2)), which is (42 * 0.7791) - (40 * e(-0.10 * 0.50))*(0.7349)    ==> $4.76

P = (K * e-rt)*N(-d2) - (S0)*N(-d1), which is (40 * e(-0.10 * 0.50))*(0.2651) - (42 * (0.2209) ==> $0.81

Value of call option is $4.76

Value of put option is $0.81

Intrinsic value of in-the-money call option = current stock price - strike price = $42 - $40 = $2

time value = option value - intrinsic value = $4.76 - $2 = $2.76

Intrinsic value of out-of-money put option = 0

time value = option value - intrinsic value = $0.81 - $0 = $0.81